### v2

# Cracking the Code: Multi-domain LLM Evaluation on Real-World Professional Exams in Indonesia

Abstract
While knowledge evaluation in large language models has predominantly focused on academic subjects like math and physics, these assessments often fail to capture the practical demands of real-world professions. In this paper, we introduce IndoCareer, a dataset comprising 8,834 multiple-choice questions designed to evaluate performance in vocational and professional certification exams across various fields. With a focus on Indonesia, IndoCareer provides rich local contexts, spanning six key sectors: (1) healthcare, (2) insurance and finance, (3) creative and design, (4) tourism and hospitality, (5) education and training, and (6) law. Our comprehensive evaluation of 27 large language models shows that these models struggle particularly in fields with strong local contexts, such as insurance and finance. Additionally, while using the entire dataset, shuffling answer options generally maintains consistent evaluation results across models, but it introduces instability specifically in the insurance and finance sectors. [^1] Data can be accessed at https://huggingface.co/datasets/indolem/IndoCareer.
Cracking the Code: Multi-domain LLM Evaluation on
Real-World Professional Exams in Indonesia
Fajri Koto Department of Natural Language Processing MBZUAI, Abu Dhabi, UAE fajri.koto@mbzuai.ac.ae
1 Introduction
The evaluation of large language models (LLMs) has shifted from traditional natural language processing (NLP) tasks Mikheev et al. (1999); Straka and Straková (2017) to more complex, knowledge-intensive, and reasoning-based challenges. One of the key datasets used to assess these abilities is the massive multitask language understanding (MMLU) Hendrycks et al. (2021). Initially introduced in English, MMLU datasets have also been developed in other languages, including Indonesian Koto et al. (2023), Chinese Li et al. (2024), and Arabic Koto et al. (2024a). These datasets consist of school exam questions across various subjects and education levels, tailored to local curricula. [^2] The English MMLU is based on the U.S. curriculum, while the Indonesian MMLU follows the Indonesian curriculum. However, they primarily focus on academic subjects, often overlooking vocational and professional expertise, which are more relevant to real-world applications.
Figure 1: Distribution of professions in IndoCareer.
Due to the recent widespread adoption of LLMs across various domains, including health Zhang et al. (2024), education Weijers et al. (2024); Srivatsa and Kochmar (2024), and finance Lee and Soon (2024), evaluating a model’s knowledge across professional fields has become crucial. For instance, in healthcare, the model must adhere to ethical standards Gundersen and Bærøe (2022) and possess expertise in prevalent regional diseases. We should not trust AI-based health recommendations from models that have not passed a competency exam. Similarly, in education, the model needs to understand and align with local government teaching guidelines. Despite the importance of certification exams in professional fields, such exams have been largely excluded from prior work Koto et al. (2023).
In this paper, we introduce IndoCareer, a dataset comprising 8,834 multiple-choice questions collected from various Indonesian competency exams, certification exams, and vocational school exams. Our focus on Indonesian addresses the limitations of prior work Koto et al. (2023) and aims to enrich language diversity and local context nuances in NLP datasets, which are predominantly English-centric Liu et al. (2024). Figure 1 shows the distribution of IndoCareer, which covers 22 different professions across 6 categories: (1) healthcare, (2) insurance and finance, (3) creative and design, (4) tourism and hospitality, (5) education and training, and (6) law. Additionally, we demonstrate that IndoCareer is generally robust to option shuffling Zhou et al. (2024) when using the entire dataset, but it specifically introduces instability in insurance and finance professions.
2 Related Work
Indonesian Language Models
IndoBERT Koto et al. (2020); Wilie et al. (2020), IndoBERTweet Koto et al. (2021), IndoGPT Cahyawijaya et al. (2021), and IndoBART Cahyawijaya et al. (2021) are among the earliest transformer-based language models developed from scratch for Indonesian. These models have been widely adopted by industry and academia across various applications. For models exceeding 1 billion parameters, no foundational models have been pre-trained exclusively on Indonesian text. Instead, research has focused on adapting multilingual models through fine-tuning techniques. Notable examples include Bactrian-X Li et al. (2023), which employs LoRA Hu et al. (2022) for fine-tuning LLama-1 Touvron et al. (2023a), and Merak Ichsan (2023), Cendol Cahyawijaya et al. (2024), and Komodo Owen et al. (2024), which are fine-tuned adaptations of LLama-2 Touvron et al. (2023b). Despite growing interest in deploying Indonesian LLMs across various domains and job sectors, there remains a lack of suitable benchmarks tailored to evaluate their performance. To address this gap, we introduce IndoCareer.
Benchmarks for Evaluating Language Models
NusaCrowd Cahyawijaya et al. (2023) represents a significant effort to consolidate scattered datasets for Indonesian NLP. While most high-quality datasets focus on classical NLP tasks such as sentiment analysis, summarization, and text classification, benchmarks for knowledge-intensive and reasoning tasks have been notably limited until very recently. The introduction of IndoMMLU Koto et al. (2023), COPAL-ID Wibowo et al. (2024), and IndoCulture Koto et al. (2024b) marks a step forward in this direction. COPAL-ID and IndoCulture focus on cultural commonsense reasoning, while IndoMMLU evaluates exam questions across different education levels in Indonesia, from primary to high school.
Despite recent advancements, a significant gap remains in evaluating LLMs on professional tasks in the Indonesian context, as IndoMMLU does not include questions from professional exams. This limitation is not unique to Indonesia; professional exam coverage is also limited in similar benchmarks for other languages. For example, English MMLU Hendrycks et al. (2021) and Chinese MMLU Li et al. (2024) include professional exam questions in only 20% of their datasets, while Arabic MMLU Koto et al. (2024a) has an even lower coverage of just 4%.
As LLMs are increasingly applied across various domains Zhang et al. (2024); Lee and Soon (2024), there is a pressing need for a benchmark that evaluates their readiness for professional job sectors. IndoCareer addresses this gap, offering a comprehensive benchmark of professional exams spanning 22 professions, making it the first of its kind in Indonesia.
3 IndoCareer
Figure 2: Example of questions in IndoCareer. The English translation is only for illustrative purposes.
IndoCareer comprises 8,834 multiple-choice questions compiled from Indonesian competency exams, certification exams, and vocational school exams across 22 professions. In Indonesia, competency exams are commonly required in healthcare professions by the government. Certification exams, on the other hand, focus on specific skills within a profession, such as tax accounting in finance. At the high school level, vocational schools offer specialized training in areas like tourism, culinary arts, and fashion design. In Figure 1, Table 1, and Table 2, we present detailed statistics for the 22 professions covered in IndoCareer. In this dataset, we exclude engineering-related professions, as their certification exams are generally conducted in English.
Data Construction
We manually collected exam questions from publicly available sources across 22 professions. A majority (78%) of the questions were sourced from Scribd, [^3] https://www.scribd.com/ a document-sharing platform, while the remaining were obtained from local government websites [^4] For example: https://badanbahasa.kemdikbud.go.id and shared Google Drive folders. We ensured that all collected questions were relevant to their respective professions and suitable for distribution for research purposes. Importantly, 99% of the exam questions were retrieved from file formats, such as PDFs and Word documents, rather than directly from web pages, minimizing the risk of overlap with training data used by LLMs.
To extract the questions and answers, we hired three professional teachers with Bachelor’s degrees in Education for a one-month period. Their task focused exclusively on text-based questions, excluding any questions containing images (see Figure 2 for examples). Each worker was responsible for extracting approximately 3,000 questions. To ensure ethical practices, they were compensated above the minimum wage in Indonesia, with the total workload equivalent to five full-time workdays.
Quality Control
We ensure the high quality of our dataset through a rigorous and multi-step quality control process. Although we employ “expert” workers who are native Indonesian speakers with at least a Bachelor’s degree, additional measures are implemented to maintain and verify quality. First, all data sources are manually checked and validated by the author before being distributed to the workers. Workers also participate in a 1-hour workshop prior to data collection, ensuring they fully understand the guidelines and the expected data standards.
After the workers complete their tasks, we apply automated filtering to eliminate repetitive questions and entries without answer keys. To further validate the dataset, we conducted a manual review of 300 randomly selected samples (3.3% of the dataset), performed by the authors of this paper. During this review, we verified the accuracy of the questions, answer options, and answer keys. The manual review achieved an accuracy rate of 99%, demonstrating the dataset’s reliability and representing the highest meaningfully achievable score for IndoCareer.
Field Professions Exam Type #Q
Healthcare Medical Doctor Competency Exam 805
Pharmacist Competency Exam 598
Midwifery Competency Exam 680
Nurse Competency Exam 497
Clinical Psychology Other 95
Insurance & Finance Life Insurance Certification Exam 476
Sharia Life Insurance Certification Exam 558
CFP Certification Exam 96
CPA Certification Exam 663
CPMA Certification Exam 169
CITA Certification Exam 253
Risk Management Certification Exam 194
Tourism & Hospitality Tourism Vocational School 222
Hospitality Vocational School 367
Culinary Art Vocational School 294
Creative & Design Graphic Design Vocational School 423
Fashion Design Vocational School 267
Broadcasting Vocational School 422
Law Advocate Certification Exam 591
Police Other 280
Education & Training Teacher Competency Test Certification Exam 538
Office Administration Vocational School 346
Table 1: Number of questions in IndoCareer across different professions. CFP stands for Certified Financial Planner, CPA stands for Certified Public Accountant, CPMA stands for Certified Professional Management Accountant, and CITA stands for Certified Indonesian Tax Accountant.
Field # Questions # Chars
Question Answer
Healthcare 2675 277.3 95.9
Insurance and Finance 2409 156.3 165.3
Tourism and Hospitality 883 99.8 96.2
Creative and Design 1112 101.0 100.5
Law 871 130.7 141.2
Education and Training 884 159.5 165.9
Table 2: Average question and answer length (in characters) for each profession fields.
Model (#parameters) Healthcare Insurance Tourism & Law Creative Education Average
& Finance Hospitality & Design & Training
Random 20.6 25.8 20.0 24.1 20.1 22.8 22.5
[…]
BLOOMZ (1.7B) 28.2 34.7 40.2 32.6 39.0 35.9 33.7
BLOOMZ (3B) 29.8 39.2 42.2 37.3 44.2 40.8 37.3
BLOOMZ (7B) 32.9 41.7 47.1 40.3 48.9 45.1 40.7
[…]
Gemma-2 (2B) 35.7 51.0 55.5 44.4 55.0 52.1 46.8
Gemma-2 (9B) 54.3 62.2 68.0 56.9 68.1 60.8 60.5
Gemma-2 (27B) 58.3 64.2 71.7 60.2 71.7 62.6 63.5
Aya-23 (8B) 37.0 46.1 51.7 44.3 51.7 47.5 44.6
Aya-23 (35B) 43.9 52.9 59.0 50.4 61.8 53.3 51.7
LLaMA-3.1 (8B) 35.9 46.7 51.9 41.2 53.0 45.3 44.1
[…]
LLaMA-3.1 (70B) 61.4 65.0 69.4 64.0 72.3 61.4 64.8
[…]
IndoGPT (117M) 21.5 26.6 24.5 23.2 18.1 23.6 23.2
Merak (7B) 37.2 45.6 49.7 43.8 50.8 46.9 44.1
SeaLLM (7B) 41.1 54.7 56.0 44.7 61.3 50.8 50.1
SEA-LION (7B) 19.2 28.9 20.0 27.6 20.9 27.3 23.8
Komodo (7B) 25.5 29.7 27.4 30.5 29.8 31.8 28.5
[…]
Table 3: Zero-shot LLM performance (accuracy), combined across professional fields. “Average” means the average across all questions in IndoCareer.
Data Statistics
Table 1 summarizes the distribution of questions in IndoCareer across 22 professions, organized into six main fields: Healthcare, Insurance & Finance, Tourism & Hospitality, Creative & Design, Law, and Education & Training. Each profession corresponds to specific exam types, including competency exams, certification exams, vocational school exams, and others. Healthcare encompasses five professions, such as Medical Doctor and Pharmacist, contributing a total of 2,675 questions. Insurance & Finance, the largest category with seven professions, includes fields like Life Insurance, Certified Public Accountant (CPA), and Risk Management, with 2,409 questions. Tourism & Hospitality covers three professions—Tourism, Hospitality, and Culinary Art—comprising 883 questions, while Creative & Design features 1,112 questions. The Law field includes Advocate and Police exams, with a total of 871 questions, while Education & Training, with Teacher Competency Tests and Office Administration, adds another 884 questions.
According to Table 2, healthcare questions are the longest, averaging 2 to 3 times the length of those in tourism and hospitality, and creative and design. The number of multiple-choice options is generally consistent across professional fields, averaging 4 options. However, the total character count of the options varies, with insurance and finance, and education and training having the longest options, exceeding 160 characters.
Additionally, we manually examined 300 random samples to assess whether answering the questions required local context. [^5] The 300 random samples are the same as those used for the manual review. Given the 99% accuracy rate from the initial review, we included an additional 1% of randomly selected correct samples for the local context assessment. Our analysis revealed that 34% of the questions incorporated Indonesian local context, with a notable concentration in the fields of insurance and finance, tourism and hospitality, and law.
4 Experiments
Pezeshkpour and Hruschka (2024); Zhou et al. (2024) demonstrated that LLMs are highly sensitive to the order of options in multiple-choice questions. To ensure a more robust evaluation, we report the average performance across three evaluations for each model: one using the original order of options and two with the options shuffled. [^6] For reproducibility, we also release two versions of IndoCareer with shuffled options, available at https://huggingface.co/datasets/indolem/IndoCareer. We evaluated one closed-source model (GPT-4o) and 26 open-weight LLMs, comprising 18 multilingual models (BLOOMZ Muennighoff et al. (2022), mT0 Muennighoff et al. (2022), Gemma-2 Team et al. (2024), Aya-23 Üstün et al. (2024), LLaMA3.1777https://github.com/meta-llama/llama3) and 8 Indonesian-centric models (IndoGPT Cahyawijaya et al. (2021), Bactrian-ID Li et al. (2023), Merak Ichsan (2023), Komodo Owen et al. (2024), SeaLLM Nguyen et al. (2023), SEA-LION Singapore (2023), and Cendol Cahyawijaya et al. (2024)). Details for each model can be found in the Appendix.
Our focus is on zero-shot experiments using the Indonesian prompt: Ini adalah soal [subject] untuk [exam type]. Pilihlah salah satu jawaban yang dianggap benar!. [^8] The English translation is ”This is a [subject] question for [exam type]. Please choose the correct answer!” For evaluation, we use the LM-Harness package Gao et al. (2024), selecting the answer based on the highest probability of the first token (i.e., A, B, C, D) in the generated output. Specifically, for GPT-4o, we used the gpt-4o model from OpenAI, [^9] https://openai.com/ selecting the answer based on the first letter generated in the output. [^10] For GPT-4o, we slightly adjusted the prompt, instructing the model to output only one of the options as the answer.
4.1 Results
Table 3 summarizes the zero-shot performance of various large language models (LLMs) across professional fields in IndoCareer, highlighting significant differences in their ability to handle Indonesian professional exams. GPT-4o and LLaMA-3.1 (70B) emerge as the top-performing models, with GPT-4o achieving the highest overall accuracy at 72.3%, followed closely by LLaMA-3.1 (70B) with 68.5%. This 4-point gap demonstrates GPT-4o’s superior capability in handling complex tasks across diverse professions. In contrast, other multilingual models show significantly lower accuracy, ranging between 38.0% and 60.0%, indicating their struggles with Indonesian-specific professional exams.
Indonesian-centric models, including SEA-LION, Komodo, and Cendol, underperform dramatically, with results close to random guessing in some fields. These findings suggest that existing Indonesian-centric models are not yet optimized for professional exam tasks, limiting their utility in practical applications. Notably, the SEA-LION (7B) and Komodo (7B) models achieve only 23.8% and 28.5% average accuracy, respectively, underscoring the gap between local adaptations and the more capable multilingual models.
Healthcare stands out as the most challenging professional field, with an average performance across all models at only 37.2%. [^11] This figure is calculated by averaging all values in the Healthcare column of Table 3. This poor performance underscores the limitations of current off-the-shelf LLMs as reliable health advisors in the Indonesian context. These findings highlight the critical need for robust model adaptations and fine-tuning specifically tailored to Indonesian professional tasks to enhance performance and to ensure applicability in high-stakes domains such as healthcare.
4.2 Analysis
Shuffling the multiple-choice options leads to unstable results in insurance and finance.
Table 4 lists the top 10 professions with the highest standard deviation ($\sigma$) in performance across three evaluation runs. While the standard deviations are relatively low, ranging from 1.5 to 3.0, they indicate minor instabilities in model predictions when the multiple-choice options are shuffled. For certain professions, such as Certified Financial Planner, Certified Indonesian Tax Accountant, and Certified Professional Management Accountant, the average rank correlation ($\tau$) drops below 0.9, indicating reduced consistency in model performance across evaluation runs. Although their deviations are not severe, they highlight areas where models are less robust to option shuffling, particularly in domains requiring nuanced reasoning. Across the entire dataset, however, the rank correlation remains high, with an average of 0.98. This indicates that while minor instabilities exist at the profession level, the overall dataset maintains stable performance.
Profession $\sigma\downarrow$ $\tau\uparrow$
Clinical Psychology 3.00 0.93
Cert. Financial Planner 2.91 0.68
Cert. Professional Management Accountant 2.00 0.90
Fashion Design 1.98 0.93
Advocate 1.96 0.91
Police 1.86 0.95
Cert. Indo. Tax Accountant 1.81 0.85
Sharia Life Insurance 1.81 0.97
Risk Management 1.76 0.97
Tourism 1.63 0.96
All 1.57 0.98
Table 4: Top 10 professions with the highest standard deviation ($\sigma$). $\tau$ represents the average rank correlation across three runs. The red cells are the three worse score. The scores are based on evaluations across 27 models.
Figure 3: Top 5 and bottom 5 professions based on the model’s accuracy disparity relative to the passing score.
LLMs perform well in life insurance certification but struggle with finance-related certifications.
Figure 3 illustrates the performance of LLMs across the top 5 and bottom 5 professions in terms of accuracy relative to the passing scores. The passing scores for each exam, represented by red horizontal lines, were sourced from publicly available information. The figure highlights that while GPT-4o, LLaMA-3.1 (70B), and Gemma-2 (27B) achieve passing scores for professions such as life insurance, sharia life insurance, graphic design, midwifery, and teacher competency, they fall significantly short for finance-related certifications.
None of the models evaluated pass the exams for Certified Financial Planner (CFP), Certified Professional Management Accountant (CPMA), Certified Public Accountant (CPA), fashion design, or office administration. Notably, GPT-4o, the best-performing model overall, falls over 20 points below the passing score for CFP, emphasizing the difficulty of finance-related tasks. The results suggest that finance-related certifications, which often require domain-specific reasoning and detailed calculations, remain a challenge for current LLMs. On the other hand, professions with more straightforward knowledge requirements, such as life insurance or midwifery, align better with the strengths of existing LLMs. These findings highlight the need for targeted fine-tuning and adaptation to improve performance in specialized and calculation-heavy fields like finance.
Questions with local context and numerical analysis pose greater challenges.
We conducted an error analysis on the best-performing open-weight model, LLaMA-3.1 (70B), by examining 100 incorrectly predicted samples and 100 correctly predicted samples for comparison. These samples were drawn from the original questions, without applying option shuffling. The analysis showed that questions with Indonesian local context were more common among the incorrectly predicted samples, with 50% of the incorrect predictions containing local context, compared to only 22% among the correct predictions. Considering that IndoCareer contains 34% local context overall, as discussed in Section 3, this suggests that questions incorporating local context are particularly challenging for language models. This finding aligns with prior research Koto et al. (2024b), indicating that questions grounded in local context often introduce cultural or situational nuances not well-captured in the models’ pretraining data.
In addition to local context, questions involving numerical analysis also posed significant challenges for LLaMA-3.1 (70B). Among the incorrectly predicted samples, 43 required numerical reasoning, compared to only 29 among the correctly predicted ones. Numerical questions often involve calculations or logical reasoning steps, which many LLMs are not explicitly optimized to handle. These results reveal two key areas where model performance could be improved: understanding and addressing culturally specific content and enhancing their capabilities for numerical reasoning.
5 Conclusion
We introduce IndoCareer as the most comprehensive dataset of professional exams across various job sectors in Indonesia. The dataset encompasses 22 professions, categorized into healthcare, insurance and finance, creative and design, tourism and hospitality, education and training, and law. Evaluations across different LLMs show that most off-the-shelf models demonstrate vocational and professional expertise below the passing scores. We believe IndoCareer will be valuable in supporting LLM adaptation for various job sectors in Indonesia.
Limitations
There are three main limitations to our work: (1) IndoCareer excludes multimodal data such as tables, audio, images, and videos. Including these would make the benchmark more comprehensive and reflective of real-world scenarios. However, since our focus is on LLM evaluation, we only include text-based questions; (2) Engineering-related professions are excluded from IndoCareer because the language used in these exams is primarily English, while our focus is on the Indonesian language; (3) The evaluation is limited to multiple-choice questions and does not include text generation tasks. We follow prior work in using the multiple-choice format as an initial step to address the lack of professional and vocational exam benchmarks in Indonesian.
Ethical Considerations
IndoCareer is released under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License [^12] https://creativecommons.org/licenses/by-nc-sa/4.0/ and is intended solely for academic research. The questions included in IndoCareer are sourced from publicly available materials. We collected these questions in compliance with Indonesian Copyright Law No. 28 of 2014, specifically Article 44. This article states that the use, reproduction, and/or modification of works or related rights, in whole or in part, is not considered copyright infringement, provided the source is properly cited and the purpose is for education or research. [^13] https://wipolex-res.wipo.int/edocs/lexdocs/laws/en/id/id064en.pdf
Appendix A Models
Models (#parameters) Source
BLOOMZ (560M) bigscience/bloomz-560m
BLOOMZ (1.1B) bigscience/bloomz-1b1
BLOOMZ (1.7B) bigscience/bloomz-1b7
BLOOMZ (3B) bigscience/bloomz-3b
BLOOMZ (7.1B) bigscience/bloomz-7b1
[…]
Gemma-2 (9B) google/gemma-2-9b-it
Gemma-2 (27B) google/gemma-2-27b-it
[…]
Aya-23 (35B) CohereForAI/aya-23-35B
[…]
LLaMA3.1-Instruct (8B) meta-llama/Meta-Llama-3.1-8B-Instruct
LLaMA3.1 (70B) meta-llama/Meta-Llama-3.1-70B
LLaMA3.1-chat (70B) meta-llama/Meta-Llama-3.1-70B-Instruct
[…]
IndoBART (132M) indobenchmark/indobart-v2
IndoGPT (117M) indobenchmark/indogpt
Merak (7B) Ichsan2895/Merak-7B-v5-PROTOTYPE1
SeaLLM (7B) SeaLLMs/SeaLLMs-v3-7B-Chat
SEA-LION (7B) aisingapore/sea-lion-7b
Komodo (7B) Yellow-AI-NLP/komodo-7b-base
[…]
Table 5: With the exception of GPT-4o, all the models used in this study were sourced from Huggingface Wolf et al. (2020).
Appendix B Full Results
Table 6 presents the accuracy of each model across various professions. The passing scores for each exam were sourced from publicly available information. We found that GPT-4o passes most of the exams, with the exceptions being Certified Financial Planner, Certified Public Accountant, Certified Professional Management Accountant (CPMA), and Office Administration. LLaMA-3.1 and Gemma-2 also pass some Indonesian exams, but no Indonesian-centric model has yet passed the professional and vocational exams in Indonesia.
Profession P.Score BLOOMZ mT0 Aya-23 Gemma-2 LLaMA3.1 Merak SeaLLM SEA-LION Komodo Cendol GPT-4o
Healthcare
Medical Doctor 66.0 33.4 27.3 45.1 61.6 70.9 39.6 42.9 19.9 24.1 23.1 74.8
Pharmacist 57.0 30.0 24.4 41.2 55.8 62.9 36.2 40.3 19.4 23.0 23.0 64.5
Midwifery 40.0 29.4 30.9 42.2 52.2 57.9 33.1 37.4 20.5 22.8 22.2 63.3
Nurse 60.0 35.3 33.9 45.5 58.4 60.8 36.2 44.7 23.8 25.6 25.7 66.4
Clinical Psycology 70.0 53.3 50.7 62.7 73.6 71.4 53.3 62.7 26.4 33.7 34.1 74.3
Insurance & Finance
Life Insurance 70.0 48.7 48.1 61.3 71.7 76.5 49.0 59.3 27.3 35.0 30.2 77.3
Sharia Life Insurance 70.0 45.9 47.1 61.6 77.8 81.7 51.7 64.1 32.2 30.0 31.8 85.3
Cert. Financial Planner 70.0 29.4 27.6 36.9 41.9 49.5 25.4 38.0 24.0 22.6 22.2 49.8
Cert. Public Accountant 75.0 37.9 36.7 44.7 58.4 61.7 42.8 48.2 25.9 27.0 26.6 67.8
Cert. Indo. Tax Accountant 60.0 37.5 39.1 41.7 47.3 50.4 34.3 45.5 32.3 33.2 31.5 60.7
CPMA 75.0 32.7 28.7 40.6 55.2 62.0 38.6 40.8 26.9 25.9 24.1 61.4
Risk Management 70.0 41.7 37.9 51.1 64.0 67.2 45.0 53.2 26.9 29.3 32.1 70.9
Tourism & Hospitality
Tourism 70.0 51.3 53.6 58.1 72.6 74.3 45.8 58.8 21.8 30.1 24.2 76.6
Hospitality 70.0 43.0 43.4 54.8 67.2 69.0 47.6 55.4 23.4 25.1 26.4 71.7
Culinary Art 70.0 47.0 45.2 62.0 73.5 76.7 50.1 60.4 22.8 28.3 22.2 79.5
Creative & Design
Fashion Design 70.0 34.8 35.2 47.1 59.0 62.4 36.4 49.1 20.6 25.3 21.7 66.2
Graphic Design 70.0 52.9 53.8 65.3 76.7 80.6 54.8 65.0 23.6 29.3 28.0 82.2
Broadcasting 70.0 51.6 49.2 63.9 75.3 77.3 54.7 63.9 23.4 30.7 25.0 79.8
Law
Advocate 70.0 34.6 39.9 47.1 59.9 68.7 36.7 41.9 26.4 27.3 26.4 72.9
Police 60.0 44.2 37.4 47.7 56.2 64.0 45.4 47.5 21.7 27.0 26.5 67.6
Education & Training
Teacher Competency 55.0 46.6 44.1 53.4 61.5 62.1 45.5 48.8 27.5 29.3 27.7 65.6
Office Administration 70.0 38.9 42.1 52.1 64.0 68.2 42.0 50.0 24.2 27.7 23.0 69.9
[…]